OpenStudy (anonymous):

simplify this quadratic further i want to verify 6(3x-2)-(3x-2)^2

OpenStudy (anonymous):

(3x-2)(6-3x+2) =(3x-2)(8-3x)

OpenStudy (anonymous):

well that was also my answer this book has errors in answers

OpenStudy (anonymous):

thank u atjari

OpenStudy (anonymous):

yw.

OpenStudy (anonymous):

atjari expand that and equate dy/dx=0 and tell me ur value for x.U can just tell me ur value no harsh

OpenStudy (anonymous):

tell me ur x value

OpenStudy (anonymous):

24x-16-9x^2+6x =30x-16-9x^2 dy/dx=30-18x 30-18x=0 18x=30 x=5/3

OpenStudy (anonymous):

sure thats gud thanks once again

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

Atjari may u do Q9 iii) please

OpenStudy (anonymous):

So sorry dr. I forget the theory of this. I remember only a very little portion of calculas. I lost touch with this for more than a year.

OpenStudy (anonymous):

thank u atjari :)

OpenStudy (anonymous):

I didn't help u nw. So I dnt thnk I deserve dat thanx. Anywy u r welcum.

OpenStudy (anonymous):

its better than no reply

OpenStudy (anonymous):

Ha k

OpenStudy (anonymous):

\[6(3x-2)-(3x-2)^{2}\] is this the equation?